A sample of pure gold has a density of 19.31 g/mL. If a cube of aluminum has a length of 5.00 cm on each side, how much mass in grams will it have?
Answer #1

From the definition of Density, d = mass / volume = m/V

Given that the length of the cube, l = 5.0 cm. But, for a cube, lenght = height =width = 5.0 cm

Volume of the cube= l^3 = 5^3 cm^3 = 125 cm^3 = 125 mL.

Density of the pure gold = m/125 = 19.31 g/mL

Therefore, mass of the pure gold = 2413.75 grams